if the mother has 48 years and her daughter has 16 before how many years was the daughter 5 times younger than her mother?

Found 2 solutions by stanbon, ankor@dixie-net.com:
Answer by stanbon(75887)

if the mother has 48 years and her daughter has 16 before how many years was the daughter 5 times younger than her mother?
---------
Let the answer be "x".
--
Equation:
48+x = 5(16+x)
------
48+x = 80+5x
4x = -32
----
x = -8
------
Mother would then be 48-8 = 40 yrs
Daughter would then be 16- 8 = 8
--------------
Cheers,
Stan H.

Answer by ankor@dixie-net.com(22740)

if the mother has 48 years and her daughter has 16 before how many years was the daughter 5 times younger than her mother?
:
let y = no. of years for this to be true
48 - y = 5(16-y)
48 - y = 80 - 5y
-y + 5y = 80 - 48
4y = 32
y = 32/4
y = 8 yrs
:
:
Check this, subtract 8 from both ages, 40 and 8, 5 * 8 = 40